# Leveraging Sparse Annotations for Leukemia Diagnosis on the Large Leukemia Dataset

### v1

@@ 3.3 LeukemiaAttri dataset
To overcome the above-mentioned dataset limitations and provide a benchmark for training deep learning models, we present a large-scale dataset called LeukemiaAttri dataset. The LeukemiaAttri dataset comprises 14 types of WBCs. The details of each type in a subset are presented in Table 3, where the “None” category includes artifacts, bare cells, and cells that were challenging for hematologists to identify. In total, the data set contains 28.9K images (2.4K images $\times$ 2 microscopes $\times$ 3 resolutions $\times$ 2 camera types), representing 12 subsets of images captured using low- and high-cost microscopes at three magnification levels. The detail of the LeukemiaAttri dataset is explained in Table 2.
@@
Sr. # Dataset Microscope Resolution Camera Image WBC Artifacts
Name Count Count
@@
Sr. # Type Count Count Sr. # Type Count Count
of cell in A in B of cell in A in B
@@ 7 Abnormal promyelocyte 276 298 14 None 2560 2375
Figure 3: Attribute classes distribution of a subset from LeukemiaAttri and Sparse leukemiaAttri datasets. The distribution shows the class imbalance in the ”Nuclear shape and Cytoplasmic vacuoles is much higher than the others.
@@ 3.3.1 Morphological attributes:
To improve prognostic analysis, hematologists identified 14 distinct types of WBCs. The morphological characteristics of several types of WBC are illustrated in Figure 2 (c). For annotating the WBC types and their morphological features, hematologists meticulously reviewed all subsets of the captured images and selected the subset with the best-quality images that offered clear structural details. This selected subset (HCM_100x_C2) , which was captured using the High-Cost Microscope (HCM) at 100x magnification with the HD1500T camera (C2) was used for the annotation process.
To ensure accuracy, two hematologists collaborated in annotating the cells. Figure 2 (c) shows examples of certain WBC types and their morphological attributes, such as monoblasts, monocytes, and myelocytes, (A, B, C in the Figure 2(c)) which exhibit similar morphological features due to their shared myeloid lineage. However, variations are observed, especially in the presence of cytoplasmic vacuoles. On the other hand, lymphoblasts (Figure 2(c): D), which belong to the lymphoid lineage, exhibit distinct morphological characteristics compared to the myeloid lineage. Once detailed annotations of WBC types and their attributes were obtained at 100x magnification using the HCM, we transferred these annotations to images captured at different resolutions and with the LCM using homography techniques [40, 41]. The transferred annotations were manually reviewed, and any missing or incorrectly localized annotations were re-annotated. The specific counts of WBC types and their corresponding attributes for the source subset are provided in Table 3
@@ 3.4 Sparse LeukemiaAttri dataset
To ensure the accuracy of morphological attributes, the LeukemiaAttri dataset was collected at 100x magnification and then the corresponding patches from 40x and 10x were extracted while ignoring the complete FOV of 40x and 10x. To make efficient use of available data and to ensure practical efficacy (since 40x is the gold standard for Leukemia), in the Sparse LeukemiaAttri dataset, the complete FOV of 40x should be considered. To address the above limitation of the LeukemiaAttri dataset, we extend the $HCM\_40X\_C2$ (see Table 2) subset to include a complete FoV. This new dataset is split up into a train set (1546 images) and a test-set (185 images). For the training set, the annotations are borrowed from the LeukemiaAttri dataset and mapped to 40x images, thus resulting in the sparse annotations and named as Sparse-LeukemiaAttri. The ratio of the labeled patch to the entire field of view (FoV) at 40x, in terms of size, is 1:5. However, for the test set (curated at 40x resolution), each patch of the image is annotated at 100x (for better annotation quality) resolution and these annotations are transferred to a complete FOV of 40x test set. Examples of train and test subsets are shown in Figure 1. Collectively, with a new extension, we have 13 subsets now as shown in Table 2.
Our Sparse-LeukemiaAttri dataset is a specialized case of sparsely supervised and semi-supervised object detection datasets. Unlike traditional sparsely supervised datasets where annotations are randomly missed, each of our images contains a fully labeled patch. Similarly, while semi-supervised approaches typically involve fully labeled and fully unlabeled sets, our dataset consists of partially annotated patches within each image, rather than entirely labeled or unlabeled images.
4 Proposed Multi Task Detectors
Our goal is to detect WBC, along with their morphological attributes, from PBF images to enable a real-time, explainable diagnosis of leukemia. To avoid large computational cost, instead of separate network we extract features from the backbone of the predicted WBC and estimate the attributes of the WBC through multi-task learning. Furthermore, to adhere with the real-word scenario, where hematologists analyze the PBF on 40x for efficiency, proposed WBC detector and attribute predictor is trained on complete FOV of 40x resolution. Since data was annotated on 100x, it maps to only small portion in 40x, resulting in sparse annotation. Therefore, we design self-learning mechanism train Multi Task object detector in a sparse setting.
@@ 4.1 Preliminaries
Let $\mathcal{F}(I_{i})=\{\hat{\mathbf{z}}_{i,k}\}_{k=1}^{n_{i}}$ be the object detector that outputs $n_{i}$ predictions. Each prediction $\hat{\mathbf{z}}_{i}^{k}=(\hat{{c}}_{i,k},\hat{{o}}_{i,k},\hat{\mathbf{p}}_{i,k},\hat{\mathbf{b}}_{i,k})$ consists of objectness score $\hat{{o}}_{i,k}$, a class with maximum probability $\hat{c}_{i,k}$, predicted probability vector $\hat{\mathbf{p}}_{i,k}$, and bounding-box $\hat{\mathbf{b}}_{i,k}\in\mathbb{R}^{4}$. The confidence score, $\hat{p}_{i,k}$, for $k^{th}$ prediction is computed as: $max(\hat{\mathbf{p}}_{i,k})\times\hat{o_{i,k}}$, where $max$ returns the maximum value from the input vector. For readability, we assume that $\hat{\mathbf{b}}_{i,k}=\{x_{k},y_{k},w_{k},h_{k}\}$ and $I_{i}\in\mathbb{R}^{H\times W}$ represents the input image and $\{\mathbf{z}_{i,k}\}_{k=1}^{n_{i}}$ is the annotation of the input image. Let $\mathcal{D_{A}}=\{(I_{i},\{\mathbf{z}_{i,k}\}_{k=1}^{n_{i}})\}_{i=1}^{N_{D}}$ be the fully annotated dataset. Each annotation $\mathbf{z}_{i}^{k}=(\mathbf{b}_{i,k},\mathbf{c}_{i,k},\mathbf{m}_{i,k})$ consists of a bounding box $\mathbf{b}_{i,k}$, a corresponding class label $\mathbf{c}_{i,k}\in\{1,\dots,C\}$, and $\mathbf{m}_{i,k}$, which is the set of morphological attributes associated with the $k^{\text{th}}$ object in the $i^{\text{th}}$ image.
Our methodology is compatible with any object detector, however, we chose YOLOv5 [17] for our approach, since our experiments indicated (Table 4) it to have higher mAP in comparison with other recent SOTA object detectors including FCOS [15], Spare-Faster-RCNN [14] and transformer-based DINO[16].
@@ 4.2 AttriDet: WBC detection with attributes
The normalized bounding box $\mathbf{b}_{\text{norm}}$ is used to extract features from $S_{j}$ followed by ROIAlign [42] to equalize the dimensions and properly align the extracted feature. Let $\mathcal{R}\mathcal{O}\mathcal{I}$ denote the ROIAlign operation.
@@
In our case, dimensions of $\mathcal{F}_{S_{1},k}$ and $\mathcal{F}_{S_{2},k}$ are 24$\times$30$\times$128 and 24$\times$30$\times$256, respectively. Finally, multi-scale features are fused $\mathcal{C_{F}}$=$\mathcal{F}_{S_{1,k}}$$\oplus$ $\mathcal{F}_{S_{2,k}}$ and then fed into the AttriHead to effectively learn the relevant WBC attributes. In AttriHead, three convolutional layers are followed by max pooling layers shown in Figure 4 (a). The output is then flattened into a d-dimensional vector and passed through a fully connected (FC) network with three layers configured as (1024, 256, 6) units. The training of AttriHead is done using multilabel binary class loss via asymmetric loss [43]. Once trained, the AttriHead predicts the six-dimensional vectors ^$\hat{\mathcal{M}}^{1\times 6}$, where each dimension represents the state of the attribute. The details of the morphological attributes are provided in Section 3.3.1. For the MTL, in addition to the AttriHead, we also train the YOLOv5 for WBC detection, utilizing the following loss functions with the standard settings of the detector:
@@
where $l$ denotes the output prediction layers, $\mathcal{L}_{total}$ represents the final loss of the object detector, which is composed of several components: $\mathcal{L}_{\text{cls}}$ represents the object classification loss, $\mathcal{L}_{\text{bbox}}$ denotes the bounding box regression loss, $\mathcal{L}_{\text{obj}}$ is the loss associated with objectness and $\mathcal{L}_{mor}$ for the attribute classification. Overall, the proposed AttriDet not only detects WBCs but also provides the detailed morphology in terms of attributes.
@@
Figure 5: Architecture of SLA-Det: The proposed framework detects sparsely annotated white blood cells (WBCs) and morphological attributes in 40× images, using a backbone for feature extraction and an AttriHead module. Labeled regions (LR) optimize predictions via labeled region loss against ground-truth labels (TL), while unlabeled regions employ pseudo-labels (PL, high confidence) and similarity pseudo labels (SL) with triplet loss to align class-based features. Training combines labeled regions, pseudo-labels, morphology-attributes, and triplet losses, enabling robust robust learning for accurate detection in sparse annotation scenarios.
@@ 4.3 SLA-Det: WBC detection with attributes using Sparse Annotation
Due to annotation cost, images in the training set of Sparse-LeukemiaAttri are sparsely annotated. Since many object detectors (e.g. Yolov5) use the IoU of the predicted bounding box against ground truth to identify the true positives and calculate the object loss, training with sparse annotations negatively impacts performance because the ground truth data restricts the model’s ability to learn effective features, causing it to struggle with differentiating between background and objects of interest. This can increase false negatives, as the model may overlook unannotated objects. To overcome these limitations sparse-annotation-based object detectors, such as SparseDet [33] and Co-mining [36], have been proposed. Usually, these methods utilize a few bounding-box annotations of the objects (with many objects present but not annotated), without much information about the background. However, the scenario on our hand is slightly more informative and requires re-designing of the solution. In our case, we have annotated regions, providing us with knowledge of where objects are present and where they are absent within those areas. Outside that region, we have no information about objects and backgrounds. For this, we propose a new sparse-annotation-based object detector of Leukemia, called Sparse Leukemia Attribute Detector (SLA-Det) is proposed.
Let $\mathcal{D_{S}}=\{(I_{i},\{\mathbf{z}_{i,k}\}_{k=1}^{n_{i}},P_{i})\}_{i=1}^{N_{D}}$ be that sparsely annotated dataset. Similar to $\mathcal{D_{A}}$, each annotation $z_{i}^{k}=(\mathbf{b}_{i,k},\mathbf{c}_{i,k},\mathbf{m}_{i,k})$ consists of bounding box $\mathbf{b}_{i,k}$, corresponding class-label $\mathbf{c}_{i,k}\in\{1,\dots,C\}$ and $\mathbf{m}_{i,k}$ is the set of morphological attributes associated with the $k^{th}$ object in the $i^{th}$ image. With each image $I_{i}$, we have associated patch ${P}_{i}$, indicating a fully annotated region. Note that this region is about 20% of the image (as shown in Figure 1). Let $\mathcal{F}(I_{i})=\{\hat{\mathbf{z}}_{i,k}\}_{k=1}^{n_{i}}$ denote the object detector that generates $n_{i}$ predictions for the image $I_{i}$. The predicted output could be separated into non-intersecting sets $\hat{\phi}_{i}$ and $\hat{\varphi}_{i}$. Where $\hat{\phi}_{i}=\{\text{set of all\leavevmode\nobreak\ }\hat{\mathbf{z}}_{i,k}\leavevmode\nobreak\ \text{that do not lie inside $P_{i}$}\}$ and $\hat{\varphi}_{i}=\{\text{set of all\leavevmode\nobreak\ }\hat{\mathbf{z}}_{i,k}\leavevmode\nobreak\ \text{that lie inside $P_{i}$}\}$.
@@ 4.3.1 Training on Annotated Regions
Handling Multiple Prediction Layers: Several object detectors ( [44, 45, 46]) generate predictions at multiple levels / layers $\{l^{v}\}_{v=1}^{N}$, to detect objects at multiple scales. In order to penalize the objectness score assigned to the background region, the location of the annotated patch has to be adjusted to these layers. We scale ${P}_{i}$ for $l^{v}$ by applying normalization (Equation 1) creating ${P}^{v}_{i,\text{norm}}$. For each level, we separate the predictions, such that $\forall_{v}\hat{\varphi}_{i}^{v}=\{\text{set of all\leavevmode\nobreak\ }\hat{\mathbf{z}}_{i,k}\text{that lie inside $P_{i}^{v}$}\}$.
Labeled region loss: To fully utilize the information in the annotated patch, the objectness loss has to be backpropagated. Since objectness is different at different layers (as each layer predicts objects of different sizes), we compute the masked objectness loss for each layer using the predication of $\hat{\varphi}_{i}^{v}$ that lies within the annotated patch. This ensures that the loss calculation focuses only on relevant regions, allowing the model to better learn from the annotated objects while ignoring irrelevant background information. Final loss for annotated region is Eq. (4):
@@ 4.3.2 Training on non-Annotated Regions
Pseudo-Label filtering: Since predictions from the not-properly trained model are noisy and might be incorrect, we design filtering criteria to select the pseudo-labels from the predictions. First, we filter out predictions, $\mathbf{z}^{*}_{i,k}$, that have predicted confidence score, ${p}^{*}_{i,k}$ below $\mathbf{t}_{0}$. This effectively means that only predictions from unlabeled regions with a high enough confidence, considering both the class probability and object score, are retained for further processing. Secondly, we remove all object predictions with an area smaller than the smallest object ($\text{Area}(\hat{b}_{i,k})>\text{Area}_{\text{min}}$) in the ground-truth in the training set. Thirdly, in the remaining set, for any prediction with confidence score $\hat{{p}}^{*}_{i,k}$ above $\mathbf{t}_{1}$, we compute entropy, allowing us to focus on the most distinguished class probabilities [47].
@@
where $\mathbf{t}_{2}$ denotes the entropy threshold. The above-selected pseudo-labels could not be used to reject the background as they only tell where the object is but not where an object could not be. Therefore, we need to mask out the object loss for the regions outside the pseudo-label. For $\hat{\mathbf{z}}_{i,k}\in D_{i}^{pseudo}$, we define the pseudo-label ${\mathbf{z}}_{i}^{*k}=(\hat{{c}}_{i,k},\hat{\mathbf{b}}_{i,k},M_{i,k})$, where $M_{i,k}$ is the mask of image size with only object location set to 1, rest is zero. Let $\phi_{i}^{*}=\{\mathbf{z}_{i}^{*k}|\hat{\mathbf{z}}_{i,k}\in D_{i}^{pseudo}$. For each level, the elements in the set are normalized per the size of the level and result in sets $\phi_{i}^{*v}$.
@@
Similarity Based Training: We hypothesize that if the classes of two objects (in our case, the object class from the labeled region versus the unlabeled region) are the same, then their features should be very similar. Conversely, if the classes are not the same, their features should not be similar. Therefore, the minimum similarity between objects of the same class must be greater than the maximum similarity between objects of different classes. Based on this hypothesis, we design a novel strategy to use $\hat{\mathbf{z}}_{i,k}\in\hat{\phi}_{i}$ which have confidence score greater than $\mathbf{t}_{0}$ indicating they might not be background but less than $\mathbf{t}_{1}$ (indicating we are not sure of the class label). Let $\hat{\mathbf{z}}^{s}_{i,k}$ be such predictions. For $\hat{\mathbf{z}}^{s}_{i,k}$, we extract the features,$\hat{f}_{i,k}$, from the initial layers of neck. Similarly, for each $j^{th}$ ground-truth bounding box in annotated patch ($P_{i}$), we extract ${f}_{i,j}$. To enhance adaptability, the triplet difference is determined by evaluating the feature maps of each filtered label against all ground truth feature maps within the batch:
@@
Similar to AttriDet, SLA-Det is built upon the YOLOv5 object detector (Fig. 5), however, the presented methodology for the AttriHead is adopted for the attribute prediction with the same setting and only for the region where the ground truth is available. Finally, to learn to predict attributes (using AttriHead) along with WBC detections, we employ the following loss.

### v2

@@ 3.3 LeukemiaAttri dataset
To overcome the above-mentioned dataset limitations and provide a benchmark for training deep learning models, we present a large-scale dataset called the LeukemiaAttri dataset. The LeukemiaAttri dataset comprises 14 types of WBCs. The details of each type in a subset are presented in Table 3, where the “None” category includes artifacts, bare cells, and cells that were challenging for hematologists to identify. In total, the data set contains 28.9K images (2.4K images $\times$ 2 microscopes $\times$ 3 resolutions $\times$ 2 camera types), representing 12 subsets of images captured using low- and high-cost microscopes at three magnification levels. The detail of the LeukemiaAttri dataset is explained in Table 2.
@@
Sr. # Dataset Name Microscope Resolution Camera Image Count WBC Count Artifacts
@@
Sr. # Type of cell Count in A Count in B Sr. # Type of cell Count in A Count in B
@@ 3.3.1 Morphological attributes:
To improve prognostic analysis, hematologists identified 14 distinct types of WBCs. The morphological characteristics of several types of WBC are illustrated in Figure 2 (c). For annotating the WBC types and their morphological features, hematologists meticulously reviewed all subsets of the captured images and selected the subset with the best-quality images that offered clear structural details. This selected subset (HCM_100x_C2), which was captured using the high-cost microscope (HCM) at 100x magnification with the HD1500T camera (C2), was used for the annotation process.
To ensure accuracy, two hematologists collaborated in annotating the cells. Figure 2 (c) shows examples of certain WBC types and their morphological attributes, such as monoblasts, monocytes, and myelocytes (A, B, C in the Figure 2(c)) which exhibit similar morphological features due to their shared myeloid lineage. However, variations are observed, especially in the presence of cytoplasmic vacuoles. On the other hand, lymphoblasts (Figure 2(c): D), which belong to the lymphoid lineage, exhibit distinct morphological characteristics compared to the myeloid lineage. Once detailed annotations of WBC types and their attributes were obtained at 100x magnification using the HCM, we transferred these annotations to images captured at different resolutions and with the LCM using homography techniques Lowe (2004); FISCHLER AND (1981). The transferred annotations were manually reviewed, and any missing or incorrectly localized annotations were re-annotated. The specific counts of WBC types and their corresponding attributes for the source subset are provided in Table 3
@@ 3.4 Sparse LeukemiaAttri dataset
To ensure the accuracy of morphological attributes, the LeukemiaAttri dataset was collected at 100x magnification, and then the corresponding patches from 40x and 10x were extracted while ignoring the complete FOV of 40x and 10x. To make efficient use of available data and to ensure practical efficacy (since 40x is the gold standard for Leukemia), in the Sparse LeukemiaAttri dataset, the complete FOV of 40x should be considered.
Figure 3: Attribute classes distribution of a subset from LeukemiaAttri and Sparse leukemiaAttri datasets. The distribution shows that the class imbalance in the Nuclear shape and Cytoplasmic vacuoles is much higher than the others.
To address the above limitation of the LeukemiaAttri dataset, we extend the HCM_40X_C2HCM\_40X\_C2 (see Table 2) subset to include a complete FoV. This new dataset is split up into a train set (1546 images) and a test-set (185 images). For the training set, the annotations are borrowed from the LeukemiaAttri dataset and mapped to 40x images, thus resulting in the sparse annotations and named as Sparse-LeukemiaAttri. The ratio of the labeled patch to the entire field of view (FoV) at 40x, in terms of size, is 1:5. However, for the test set (curated at 40x resolution), each patch of the image is annotated at 100x (for better annotation quality) resolution, and these annotations are transferred to a complete FOV of 40x test set. Examples of train and test subsets are shown in Figure 1. Collectively, with a new extension, we have 13 subsets now as shown in Table 2.
Our Sparse-LeukemiaAttri dataset is a specialized case of sparsely supervised and semi-supervised object detection datasets. Unlike traditional sparsely supervised datasets, where annotations are randomly missed, each of our images contains a fully labeled patch. Similarly, while semi-supervised approaches typically involve fully labeled and fully unlabeled sets, our dataset consists of partially annotated patches within each image, rather than entirely labeled or unlabeled images.
4 Proposed Multi-Task Detectors
Our goal is to detect WBC, along with their morphological attributes, from PBF images to enable a real-time, explainable diagnosis of leukemia. To avoid a large computational cost, instead of a separate network, we extract features from the backbone of the predicted WBC and estimate the attributes of the WBC through multi-task learning. Furthermore, to adhere with the real-world scenario, where hematologists analyze the PBF on 40x for efficiency, the proposed WBC detector, including the attribute predictor, is trained on the complete FOV of 40x resolution. Since data was annotated on 100x, it maps to only a small portion in 40x, resulting in sparse annotation. Therefore, we design a self-learning mechanism to train a multi-task object detector in a sparse setting.
@@ 4.1 Preliminaries
Let $\mathcal{F}(I_{i})=\{\hat{\mathbf{z}}_{i,k}\}_{k=1}^{n_{i}}$ be the object detector that outputs $n_{i}$ predictions. Each prediction $\hat{\mathbf{z}}_{i}^{k}=(\hat{{c}}_{i,k},\hat{{o}}_{i,k},\hat{\mathbf{p}}_{i,k},\hat{\mathbf{b}}_{i,k})$ consists of objectness score $\hat{{o}}_{i,k}$, a class with maximum probability $\hat{c}_{i,k}$, predicted probability vector $\hat{\mathbf{p}}_{i,k}$, and bounding-box $\hat{\mathbf{b}}_{i,k}\in\mathbb{R}^{4}$. The confidence score, $\hat{p}_{i,k}$, for $k^{th}$ prediction is computed as: $max(\hat{\mathbf{p}}_{i,k})\times\hat{o_{i,k}}$, where maxmax returns the maximum value from the input vector. For readability, we assume that $\hat{\mathbf{b}}_{i,k}=\{x_{k},y_{k},w_{k},h_{k}\}$ and $I_{i}\in\mathbb{R}^{H\times W}$ represents the input image and $\{\mathbf{z}_{i,k}\}_{k=1}^{n_{i}}$ is the annotation of the input image. Let $\mathcal{D_{A}}=\{(I_{i},\{\mathbf{z}_{i,k}\}_{k=1}^{n_{i}})\}_{i=1}^{N_{D}}$ be the fully annotated dataset. Each annotation $\mathbf{z}_{i}^{k}=(\mathbf{b}_{i,k},\mathbf{c}_{i,k},\mathbf{m}_{i,k})$ consists of a bounding box $\mathbf{b}_{i,k}$, a corresponding class label $\mathbf{c}_{i,k}\in\{1,\dots,C\}$, and $\mathbf{m}_{i,k}$, which is the set of morphological attributes associated with the $k^{\text{th}}$ object in the $i^{\text{th}}$ image.
Our methodology is compatible with any object detector, however, we chose YOLOv5 Ultralytics (2021) for our approach, since our experiments indicated (Table 4) it to have higher mAP in comparison with other recent SOTA object detectors including FCOS Tian et al. (2019), Spare-Faster-RCNN Sun et al. (2021) and transformer-based DINOZhang et al. (2022).
@@ 4.2 AttriDet: WBC detection with attributes
The normalized bounding box $\mathbf{b}_{\text{norm}}$ is used to extract features from $S_{j}$ followed by ROIAlign He et al. (2017) to equalize the dimensions and properly align the extracted feature. Let $\mathcal{R}\mathcal{O}\mathcal{I}$ denote the ROIAlign operation.
@@
[…]
@@
where $l$ denotes the output prediction layers, $\mathcal{L}_{total}$ represents the final loss of the object detector, which is composed of several components: $\mathcal{L}_{\text{cls}}$ represents the object classification loss, $\mathcal{L}_{\text{bbox}}$ denotes the bounding box regression loss, $\mathcal{L}_{\text{obj}}$ is the loss associated with objectness and $\mathcal{L}_{mor}$ for the attribute classification. Overall, the proposed AttriDet not only detects WBCs but also provides detailed morphology in terms of attributes.
@@
Figure 5: Architecture of SLA-Det: The proposed framework detects sparsely annotated white blood cells (WBCs) and morphological attributes in 40× images, using a backbone for feature extraction and an AttriHead module. Labeled regions (LR) optimize predictions via labeled region loss against ground-truth labels (TL), while unlabeled regions employ pseudo-labels (PL, high confidence) and similarity pseudo-labels (SL) with triplet loss to align class-based features. Training combines labeled regions, pseudo-labels, morphology-attributes, and triplet losses, enabling robust learning for accurate detection in sparse annotation scenarios.
@@ 4.3 SLA-Det: WBC detection with attributes using sparse annotation
Due to annotation cost, images in the training set of sparse-leukemiaAttri are sparsely annotated. Since many object detectors (e.g. Yolov5) use the IoU of the predicted bounding box against ground truth to identify the true positives and calculate the object loss, training with sparse annotations negatively impacts performance because the ground truth data restricts the model’s ability to learn effective features, causing it to struggle with differentiating between background and objects of interest. This can increase false negatives, as the model may overlook unannotated objects. To overcome these limitations, sparse-annotation-based object detectors, such as SparseDet Suri et al. (2023) and Co-mining Wang et al. (2021), have been proposed. Usually, these methods utilize a few bounding-box annotations of the objects (with many objects present but not annotated), without much information about the background. However, the scenario at hand is slightly more informative and requires redesigning the solution. In our case, we have annotated regions, providing us with knowledge of where objects are present and where they are absent within those areas. Outside that region, we have no information about objects and backgrounds. For this, we propose a new sparse annotation-based object detector of leukemia, called sparse leukemia attribute detector (SLA-Det) is proposed.
[…]
@@ 4.3.1 Training on Annotated Regions
Handling multiple prediction layers: Several object detectors ( Qiao et al. (2021); Lin et al. (2017); Tan et al. (2020)) generate predictions at multiple levels / layers $\{l^{v}\}_{v=1}^{N}$, to detect objects at multiple scales. In order to penalize the objectness score assigned to the background region, the location of the annotated patch has to be adjusted to these layers. We scale ${P}_{i}$ for $l^{v}$ by applying normalization (Equation 1) creating ${P}^{v}_{i,\text{norm}}$. For each level, we separate the predictions, such that $\forall_{v}\hat{\varphi}_{i}^{v}=\{\text{set of all\leavevmode\nobreak\ }\hat{\mathbf{z}}_{i,k}\text{that lie inside $P_{i}^{v}$}\}$. Labeled region loss: To fully utilize the information in the annotated patch, the objectness loss has to be backpropagated. Since objectness is different at different layers (as each layer predicts objects of different sizes), we compute the masked objectness loss for each layer using the prediction of $\hat{\varphi}_{i}^{v}$ that lies within the annotated patch. This ensures that the loss calculation focuses only on relevant regions, allowing the model to better learn from the annotated objects while ignoring irrelevant background information. Final loss for annotated region is Eq. (4):
@@ 4.3.2 Training on non-Annotated Regions
Pseudo-Label filtering: Since predictions from the not-properly trained model are noisy and might be incorrect, we design filtering criteria to select the pseudo-labels from the predictions. First, we filter out predictions, $\mathbf{z}^{*}_{i,k}$, that have predicted confidence score, ${o}^{*}_{i,k}$ below $\mathbf{t}_{0}$. This effectively means that only predictions from unlabeled regions with a high enough confidence, considering both the class probability and object score, are retained for further processing. Secondly, we remove all object predictions with an area smaller than the smallest object ($\text{Area}(\hat{b}_{i,k})>\text{Area}_{\text{min}}$) in the ground-truth in the training set. Thirdly, in the remaining set, for any prediction with confidence score $\hat{{o}}^{*}_{i,k}$ above $\mathbf{t}_{1}$, we compute entropy, allowing us to focus on the most distinguished class probabilities Grandvalet and Bengio (2004):
@@
where $\mathbf{t}_{2}$ denotes the entropy threshold. The above-selected pseudo-labels could not be used to reject the background as they only tell where the object is, but not where an object could not be. Therefore, we need to mask out the object loss for the regions outside the pseudo-label. For $\hat{\mathbf{z}}_{i,k}\in D_{i}^{pseudo}$, we define the pseudo-label ${\mathbf{z}}_{i}^{*k}=(\hat{{c}}_{i,k},\hat{\mathbf{b}}_{i,k},M_{i,k})$, where $M_{i,k}$ is the mask of image size with only object location set to 1, rest is zero. Let $\phi_{i}^{*}=\{\mathbf{z}_{i}^{*k}|\hat{\mathbf{z}}_{i,k}\in D_{i}^{pseudo}$ }. For each level, the elements in the set are normalized per the size of the level and result in sets $\phi_{i}^{*v}$.
@@
[…]
@@
Similar to AttriDet, SLA-Det is built upon the YOLOv5 object detector (Fig. 5); however, the presented methodology for the AttriHead is adopted for the attribute prediction with the same setting and only for the region where the ground truth is available. Finally, to learn to predict attributes (using AttriHead) along with WBC detections, we employ the following loss.